# Stable Flow: Vital Layers for Training-Free Image Editing

### v1

@@ 1 Introduction
Figure 3: Layer Removal. (Left) Text-to-image DiT models consist of consecutive layers connected through residual connections [33]. Each layer implements a multimodal diffusion transformer block [25] that processes a combined sequence of text and image embeddings. (Right) For each DiT layer, we performe an ablation by bypassing the layer using its residual connection. Then, we compare the generated result on the ablated model with the complete model using a perceptual similarity metric.
@@ 2 Related Work
After the emergence of text-to-image models, many works have suggested using them for various applications [7, 37, 20, 12, 75], including text-driven image editing tasks [39, 65]. SDEdit [52] addressed the image-to-image translation task adding noise and denoising with a new prompt. Blended Diffusion [8, 9] suggested performing localized image editing [38, 47, 15, 55, 62] incorporating an input mask into the diffusion process in a training-free manner, while GLIDE [54], ImagenEditor [89], and SmartBrush [93] offered fine-tuning the model on a given input mask. Other works suggested inferring the mask from the input image [21, 88] or via an additional click input [69]. Other works [34, 84, 60, 18] offered to inject information from the input image using parallel generation. While some methods [13, 45, 85] suggested fine-tuning the model per-image, a recent line of work suggested training a designated model on a large synthetic dataset for instruction-based edits [17, 97, 76]. However, none of the above methods is a training-free method that supports non-rigid editing, object adding/replacement and scene editing altogether. Concurrency, Add-it [81] offers a solution to the task of object addition using FLUX. Our method, on the other hand, supports other types of image editing (e.g., non-rigid editing, object replacement).
@@ 3.1 Measuring the Importance of DiT Layers
Figure 6: Multi-Modal Attention Distribution. Given an input image of a man, we edit it to hold an avocado by injecting the reference image in the vital layers only (left) or in the non-vital layers (right), and visualize the multimodal attention of two points: a yellow point in a region that should remain unchanged (requiring copying from the reference image), and a red point in an area targeted for editing (requiring generation based on the text prompt). As can be seen, in vital layers (left), points meant to remain unchanged show dominant attention to visual features, while points targeted for editing exhibit stronger attention to relevant text tokens (e.g., “avocado”). Conversely, non-vital layers (right) show predominantly image-based attention even in regions marked for editing. This suggests that injecting features into vital layers strikes a good multimodal attention balance between preserving source content and incorporating text-guided modifications.
@@ 3.3 Latent Nudging for Real Image Editing
However, as Figure 7(a) demonstrates, this approach proves insufficient for FLUX, resulting in corrupted image reconstructions and unintended modifications during editing. We hypothesize that the assumption $u(z_{t})\approx u(z_{t-1})$ does not hold, which causes the model to significantly alter the image during the forward process. To address this, we introduce latent nudging: multiplying the initial latent $z_{0}$ by a small scalar $\lambda=1.15$ to slightly offset it from the training distribution. While this modification is visually imperceptible (Figure 7(b)), it significantly reduces reconstruction errors and constrains edits to the intended regions. See the supplementary material for more details
@@ A.1 Method Implementation Details
As described in Section 3.1 of the main paper, we started by collecting a dataset of $K=64$ text prompts using ChatGPT [56]. We instructed it to generate text prompts describing a diverse set of objects in different environments, with the focus on one main object. Then, we sampled $K$ seeds denoted by $S$ and used them to generate $K$ corresponding images $G_{\textit{ref}}$. Next, for each layer $l$, we bypass it by taking only the residual connection values. For each bypass, we generate $K$ images using the same seed set $S$ demoted by $G_{l}$. All the images were generated using Euler sampler in 15 steps and a guidance scale of $3.5$.
@@ A.2 Baselines Implementation Details
As explained in Section 4.1 of the main paper, we compare our method against the following baselines: SDEdit [52], P2P+NTI [34, 53], Instruct-P2P [17], MagicBrush [97], and MasaCTRL [18]. We reimplement SDEdit using the FLUX.1-dev model [46], and use the official implementation for the rest of the baselines.
@@ A.3 Automatic Metrics Implementation Details
As explained in Section 4.1 of the main paper, we prepare an evaluation dataset based on the COCO [48] validation dataset. We begin by filtering the dataset automatically to include at least one prominent non-rigid body. More specifically, we filter only images containing humans or animals that at least one of them is prominent enough, but not too small, i.e., the prominent non-rigid body occupies at least 5% of the image but no more than 33%. Next, for each image, we apply various image editing tasks (non-rigid editing, object addition, object replacement, and scene editing) that take into account the prominent object from a list of different combinations, resulting in a total dataset of 3,200 samples. Examples of images from this dataset can be seen in Figure 13.
@@ A.4 User Study Details
As described in Section 4.2 of the main paper, we conducted an extensive user study using the Amazon Mechanical Turk (AMT) [6] platform, using automatically generated test examples, as explained in Section A.3. We compared all the baselines with our method using a standard two-alternative forced-choice format. The users were given full instructions, as can be seen in Figure 11. Then, for each study trial, as shown in Figure 12, users were presented with an image and an instruction “Given the following input image of a {CATEGORY} where {CATEGORY} is the COCO category of the prominent object. The users were given two editing results — one from our method and one from the baseline, and were asked the following questions:
@@ B.1 Additional Comparisons and Results
Figure 13: Baselines Qualitative Comparison on Automatic Dataset. As explained in Section 4.1 of the main paper, we compare our method against the baselines on real images extracted from the COCO [48] dataset. We find that SDEdit [52] struggles with preserving the object identities and backgrounds (e.g., bear and chicken examples). P2P+NTI  [34, 53] struggles with preserving object identities (e.g., bear and person examples) and with adding new objects (e.g., missing hat in the sheep example and missing ball in the elephant example). Instruct-P2P [17] and MagicBrush [97] struggle with non-rigid editing (e.g., person raising hand). MasaCTRL [18] struggles with preserving object identities (e.g., bear and person examples) and adding new objects (e.g., sheep and cat examples). Our method, on the other hand, is able to adhere to the editing prompt while preserving the identities.
@@
Figure 14: Ablations Qualitative Comparison on Automatic Dataset. As explained in Section 4.3 of the main paper, we compare our method against several ablation cases on real images extracted from the COCO [48] dataset. As can be seen, we found that (1) performing attention injection in all the layers or performing (3) an attention extension in all the layers encourages the model to directly copy the input image while neglecting the target prompt. In addition, (2) performing an attention extension in the non-vital layers or (4) removing the latent nudging reduces the input image similarity significantly.
@@
In Figure 13 we provide an additional qualitative comparison of our method against the baselines on real images extracted from the COCO [48] dataset, as explained in Section 4.1 in the main paper. As can be seen, SDEdit [52] struggles with preserving the object identities and backgrounds (e.g., the bear and chicken examples). P2P+NTI  [34, 53] struggles with preserving object identities (e.g., the bear and person examples) and with adding new objects (e.g., the missing hat in the sheep example and missing ball in the elephant example). Instruct-P2P [17] and MagicBrush [97] struggle with non-rigid editing (e.g., the person raising hand example). MasaCTRL [18] struggles with preserving object identities (e.g., the bear and person examples) and adding new objects (e.g., the sheep and cat examples). Our method, on the other hand, is able to adhere to the editing prompt while preserving the identities.
Next, in Figure 14, we provide a qualitative comparison of the ablated cases that are explained in, Section 4.3 in the main paper. As can be seen, we found that (1) performing attention injection in all the layers or performing (3) an attention extension in all the layers, encourages the model to directly copy the input image while neglecting the target prompt. In addition, (2) performing an attention extension in the non-vital layers or (4) removing the latent nudging reduces the input image similarity significantly.
Finally, in Figures 15 and 16, we present additional image editing results using our method.
@@ B.2 Different Perceptual Metrics
As explained in Section 3.1 of the main paper, we assess the impact of each layer by measuring the perceptual similarity between $G_{\textit{ref}}$ and $G_{\ell}$ using DINOv2 [57]. It raises the question of the importance of the specific perceptual [42] similarity metric when determining the vital layers.
To this end, we also experiment with different perceptual metrics: DINOv1 [19], CLIP [67], and LPIPS [98]. In Figures 18, 19 and 20 we plot the perceptual similarity per layer for each of these metrics. The vital layers, ordered by vitality, as defined in Equation 1 of the main paper, for each metric are:
@@
B.3 Number of Vital Layers
@@
Figure 22: Number of Vital Layers Comparison. As explained in Section B.3, we experimented with choosing a different portion of the calculated vital layer set $V$. As can be seen, when removing $80\%$ of the vital layer set, the changes are negligible. However, when removing more than that, the editing results include unintended changes, such as identity changes (e.g., the man and woman examples) and background changes (e.g., the cat and blackboard examples).
The somewhat agnostic nature of our method to the specific perceptual metric, as described in Section B.2, raises the question of the importance of the entire vital layer set $V$ to the editing task. To this end, in Figure 22 we experimented with omitting a growing number of vital layers and testing the editing results. As can be seen, when removing $80\%$ of the vital layer set, the changes are negligible. However, when removing more than that, the editing results include unintended changes, such as identity changes (e.g., man and woman examples) and background changes (e.g., cat and blackboard examples). This is consistent with the results from Section B.2 that show that the least vital layers for each perceptual metric are less important for the image editing task.
B.4 Latent Nudging Experiments
@@
B.5 Layer Bypassing Visualization
@@
As explained in Section 3.1 of the main paper, to quantify layer importance in FLUX model, we devised a systematic evaluation approach. Using ChatGPT [56], we automatically generated a set $P$ of $k=64$ diverse text prompts, and draw a set $S$ of random seeds. Each of these prompts was used to generate a reference image, yielding a set $G_{\textit{ref}}$. For each DiT layer $\ell\in\mathbb{L}$, we performed an ablation by bypassing the layer using its residual connection. This process generated a set of images $G_{\ell}$ from the same prompts and seeds.
@@
B.6 Stable Diffusion 3 Results
@@
As described in Section 3.1 of the main paper, we measured the importance of each of the layers of this model. As shown in Figure 33, we measured the effect of removing each layer from the model by calculating the perceptual similarity between the generated images with and without this layer. Lower perceptual similarity indicates significant changes in the generated images. As can be seen, removing certain layers significantly affects the generated images, while others have minimal impact.
@@
Finally, in Figure 35, we perform various editing operations using the same mechanism of injecting the reference image information into the vital layers of the model, as described in Section 3.2 of the main paper.

### v2

@@ 1 Introduction
Figure 3: Layer Removal. (Left) Text-to-image DiT models consist of consecutive layers connected through residual connections [34]. Each layer implements a multimodal diffusion transformer block [26] that processes a combined sequence of text and image embeddings. (Right) For each DiT layer, we perform an ablation by bypassing the layer using its residual connection. Then, we compare the generated result on the ablated model with the complete model using a perceptual similarity metric.
@@ 2 Related Work
After the emergence of text-to-image models, many works have suggested using them for various applications [8, 38, 21, 13, 76], including text-driven image editing tasks [40, 66]. SDEdit [53] addressed the image-to-image translation task adding noise and denoising with a new prompt. Blended Diffusion [9, 10] suggested performing localized image editing [39, 48, 16, 56, 63] incorporating an input mask into the diffusion process in a training-free manner, while GLIDE [55], ImagenEditor [89], and SmartBrush [93] offered fine-tuning the model on a given input mask. Other works suggested inferring the mask from the input image [22, 88] or via an additional click input [70]. Other works [35, 84, 61, 19] offered to inject information from the input image using parallel generation. While some methods [14, 46, 85, 100, 99] suggested fine-tuning the model per-image, a recent line of work suggested training a designated model on a large synthetic dataset for instruction-based edits [18, 97, 77]. However, none of the above methods is a training-free method that supports non-rigid editing, object adding/replacement and scene editing altogether.
@@ 3.1 Measuring the Importance of DiT Layers
Figure 6: Multi-Modal Attention Distribution. Given an input image of a man, we edit it to hold an avocado by injecting the reference image attention activations in the vital layers only (left), or in the non-vital layers (right), and visualize the multimodal attention of two points: a yellow point in a region that should remain unchanged (requiring copying from the reference image), and a red point in an area targeted for editing (requiring generation based on the text prompt). As can be seen, in vital layers (left), points meant to remain unchanged show dominant attention to visual features, while points targeted for editing exhibit stronger attention to relevant text tokens (e.g., “avocado”). Conversely, non-vital layers (right) show predominantly image-based attention even in regions marked for editing. This suggests that injecting features into vital layers strikes a good multimodal attention balance between preserving source content and incorporating text-guided modifications.
@@ 3.3 Latent Nudging for Real Image Editing
However, as Figure 7(a) demonstrates, this approach proves insufficient for FLUX, resulting in corrupted image reconstructions and unintended modifications during editing. We hypothesize that the assumption $u(z_{t})\approx u(z_{t-1})$ does not hold, which causes the model to significantly alter the image during the forward process. To address this, we introduce latent nudging: multiplying the initial latent $z_{0}$ by a small scalar $\lambda=1.15$ to slightly offset it from the training distribution. While this modification is visually imperceptible (Figure 7(b)), it significantly reduces reconstruction errors and constrains edits to the intended regions. See the supplementary material for more details and ablations.
@@ A.1 Method Implementation Details
Figure 11: Attention Injection. We adapt the self-attention injection mechanism, previously shown effective for image editing in UNet-based diffusion models, to the DiT-based FLUX architecture. Since each DiT layer processes a sequence of image and text embeddings, we propose generating both the reference image $x$ and generated image $\hat{x}$ in parallel while selectively replacing the attention keys and values that correspond to the image embeddings of $\hat{x}$ with those of $x$. This replacement is performed only within the vital layers set.
As described in Section 3.1, we started by collecting a dataset of $k=64$ text prompts using ChatGPT [57]. We instructed it to generate text prompts describing a diverse set of objects in different environments, with the focus on one main object. Then, we sampled $k$ seeds denoted by $S$ and used them to generate $k$ corresponding images $G_{\textit{ref}}$. Next, for each layer $l$, we bypass it by taking only the residual connection values. For each bypass, we generate $k$ images using the same seed set $S$ demoted by $G_{l}$. All the images were generated using Euler sampler in 15 steps and a guidance scale of $3.5$.
@@
In addition, as mentioned in Section 3.2, We adapt the self-attention injection mechanism, previously to be effective for image and video editing [91, 19] in UNet-based diffusion models, to the DiT-based FLUX architecture. Since each DiT layer processes a sequence of image and text embeddings, we propose generating both the reference image $x$ and generated image $\hat{x}$ in parallel while selectively replacing the image embeddings of $\hat{x}$ with those of $x$, but only within the vital layers set. A full visualization can be found in Figure 11.
Lastly, the variance list of the perceptual similarity of the different layers, as explained in Section 3.1, is as follows: [0.222, 0.041, 0.076, 0.08, 0.123, 0.101, 0.135, 0.124, 0.112, 0.105, 0.097, 0.12, 0.118, 0.086, 0.116, 0.067, 0.065, 0.116, 0.146, 0.065, 0.098, 0.061, 0.076, 0.077, 0.072, 0.086, 0.069, 0.067, 0.081, 0.091, 0.074, 0.062, 0.061, 0.044, 0.04, 0.054, 0.036, 0.038, 0.037, 0.04, 0.066, 0.04, 0.034, 0.044, 0.044, 0.031, 0.033, 0.036, 0.03, 0.032, 0.026, 0.026, 0.026, 0.079, 0.039, 0.037, 0.026].
@@ A.2 Baselines Implementation Details
As explained in Section 4.1, we compare our method against the following baselines: SDEdit [53], P2P+NTI [35, 54], Instruct-P2P [18], MagicBrush [97], and MasaCTRL [19]. We reimplement SDEdit using the FLUX.1-dev model [47], and use the official implementation for the rest of the baselines.
@@ A.3 Automatic Metrics Implementation Details
As explained in Section 4.1, we prepare an evaluation dataset based on the COCO [49] validation dataset. We begin by filtering the dataset automatically to include at least one prominent non-rigid body. More specifically, we filter only images containing humans or animals that at least one of them is prominent enough, but not too small, i.e., the prominent non-rigid body occupies at least 5% of the image but no more than 33%. Next, for each image, we apply various image editing tasks (non-rigid editing, object addition, object replacement, and scene editing) that take into account the prominent object from a list of different combinations, resulting in a total dataset of 3,200 samples. Examples of images from this dataset can be seen in Figure 14.
@@ A.4 User Study Details
As described in Section 4.2 we conducted an extensive user study using the Amazon Mechanical Turk (AMT) [7] platform, using automatically generated test examples, as explained in Section A.3. We compared all the baselines with our method using a standard two-alternative forced-choice format. The users were given full instructions, as can be seen in Figure 12. Then, for each study trial, as shown in Figure 13, users were presented with an image and an instruction “Given the following input image of a {CATEGORY} where {CATEGORY} is the COCO category of the prominent object. The users were given two editing results — one from our method and one from the baseline, and were asked the following questions:
@@ B.1 Additional Comparisons and Results
Figure 14: Baselines Qualitative Comparison on Automatic Dataset. As explained in Section 4.1, we compare our method against the baselines on real images extracted from the COCO [49] dataset. We find that SDEdit [53] struggles with preserving the object identities and backgrounds (e.g., bear and chicken examples). P2P+NTI  [35, 54] struggles with preserving object identities (e.g., bear and person examples) and with adding new objects (e.g., missing hat in the sheep example and missing ball in the elephant example). Instruct-P2P [18] and MagicBrush [97] struggle with non-rigid editing (e.g., person raising hand). MasaCTRL [19] struggles with preserving object identities (e.g., bear and person examples) and adding new objects (e.g., sheep and cat examples). Our method, on the other hand, is able to adhere to the editing prompt while preserving the identities.
@@
Figure 15: Ablations Qualitative Comparison on Automatic Dataset. As explained in Section 4.3, we compare our method against several ablation cases on real images extracted from the COCO [49] dataset. As can be seen, we found that (1) performing attention injection in all the layers or performing (3) an attention extension in all the layers encourages the model to directly copy the input image while neglecting the target prompt. In addition, (2) performing an attention extension in the non-vital layers or (4) removing the latent nudging reduces the input image similarity significantly.
@@
In Figure 14 we provide an additional qualitative comparison of our method against the baselines on real images extracted from the COCO [49] dataset, as explained in Section 4.1. As can be seen, SDEdit [53] struggles with preserving the object identities and backgrounds (e.g., the bear and chicken examples). P2P+NTI  [35, 54] struggles with preserving object identities (e.g., the bear and person examples) and with adding new objects (e.g., the missing hat in the sheep example and missing ball in the elephant example). Instruct-P2P [18] and MagicBrush [97] struggle with non-rigid editing (e.g., the person raising hand example). MasaCTRL [19] struggles with preserving object identities (e.g., the bear and person examples) and adding new objects (e.g., the sheep and cat examples). Our method, on the other hand, is able to adhere to the editing prompt while preserving the identities.
Next, in Figure 15, we provide a qualitative comparison of the ablated cases that are explained in, Section 4.3. As can be seen, we found that (1) performing attention injection in all the layers or performing (3) an attention extension in all the layers, encourages the model to directly copy the input image while neglecting the target prompt. In addition, (2) performing an attention extension in the non-vital layers or (4) removing the latent nudging reduces the input image similarity significantly.
Finally, in Figures 16 and 17, we present additional image editing results using our method.
@@ B.2 Different Perceptual Metrics
As explained in Section 3.1, we assess the impact of each layer by measuring the perceptual similarity between $G_{\textit{ref}}$ and $G_{\ell}$ using DINOv2 [58]. It raises the question of the importance of the specific perceptual [43] similarity metric when determining the vital layers.
To this end, we also experiment with different perceptual metrics: DINOv1 [20], CLIP [68], and LPIPS [98]. In Figures 19, 20 and 21 we plot the perceptual similarity per layer for each of these metrics. The vital layers, ordered by vitality, as defined in Equation 1, for each metric are:
@@
B.3 VLM-Based Quantiative Metric
Table 5: VLM-Based quantitative comparison. For each method, we used Phi-3.5-vision [3] VLM to compute the percentage of editing results that follow the text prompt and of the results that change only the essential parts of the image. P2P+NTI [35, 54], Instruct-P2P [18], and MasaCTRL [19] suffer from low similarity to the text prompt. SDEdit [94] and MagicBrush [97] adhere more to the text prompt, but they struggle with avoiding unintended changes.
Method Text Following ()($\uparrow$)( ↑ ) Modify only essential ()($\uparrow$)( ↑ ) SDEdit [88] 86.66% 21.66% P2P+NTI [33, 51] 68.33% 26.66% Instruct-P2P [17] 33.33% 26.66% MagicBrush [91] 88.33% 46.66% MasaCTRL [18] 33.33% 06.66% Stable Flow (ours) 83.33% 61.66%
As explained in Section 4.1, We evaluated the editing results using three widely-used CLIP-based metrics: $\text{CLIP}_{img}$, $\text{CLIP}_{txt}$, and $\text{CLIP}_{dir}$. In addition, we experimented with a VLM-based metric using the Phi-3.5-vision [3] VLM that was trained specifically for the task of multiple image comparison. For each input image $x$, editing prompt $p$, and editing result $\hat{x}$, we computed the following two metrics: (1) Text Following — we presented the VLM the edited image $\hat{x}$ and the editing prompt $p$ and asked it “Does this image correspond to the text $p$? Answer yes or no.”. (2) Modify only essential parts — we extracted the prompt instruction and presented it, along with $x$ and $\hat{x}$, to the VLM and asked it “Is the only difference between these two images the text PROMPT? Answer yes or no”. For each metric, we calculated the number of times that the VLM answered “yes”. As demonstrated in Table 5, the VLM-based metric follows the same trend as the CLIP-based metrics:P2P+NTI [35, 54], Instruct-P2P [18], and MasaCTRL [19] suffer from low similarity to the text prompt. SDEdit [94] and MagicBrush [97] adhere more to the text prompt, but they struggle with avoiding unintended changes.
B.4 Number of Vital Layers
@@
Figure 23: Number of Vital Layers Comparison. As explained in Section B.4, we experimented with choosing a different portion of the calculated vital layer set $V$. As can be seen, when removing $20\%$ of the vital layer set, the changes are negligible. However, when removing more than that, the editing results include unintended changes, such as identity changes (e.g., the man and woman examples) and background changes (e.g., the cat and blackboard examples).
The somewhat agnostic nature of our method to the specific perceptual metric, as described in Section B.2, raises the question of the importance of the entire vital layer set $V$ to the editing task. To this end, in Figure 23 we experimented with omitting a growing number of vital layers and testing the editing results. As can be seen, when removing $20\%$ of the vital layer set, the changes are negligible. However, when removing more than that, the editing results include unintended changes, such as identity changes (e.g., man and woman examples) and background changes (e.g., cat and blackboard examples). This is consistent with the results from Section B.2 that show that the least vital layers for each perceptual metric are less important for the image editing task.
B.5 Latent Nudging Experiments
@@
B.6 Layer Bypassing Visualization
@@
As explained in Section 3.1, to quantify layer importance in FLUX model, we devised a systematic evaluation approach. Using ChatGPT [57], we automatically generated a set $P$ of $k=64$ diverse text prompts, and draw a set $S$ of random seeds. Each of these prompts was used to generate a reference image, yielding a set $G_{\textit{ref}}$. For each DiT layer $\ell\in\mathbb{L}$, we performed an ablation by bypassing the layer using its residual connection. This process generated a set of images $G_{\ell}$ from the same prompts and seeds.
@@
B.7 Stable Diffusion 3 Results
@@
As described in Section 3.1, we measured the importance of each of the layers of this model. As shown in Figure 34, we measured the effect of removing each layer from the model by calculating the perceptual similarity between the generated images with and without this layer. Lower perceptual similarity indicates significant changes in the generated images. As can be seen, removing certain layers significantly affects the generated images, while others have minimal impact.
@@
Finally, in Figure 36, we perform various editing operations using the same mechanism of injecting the reference image information into the vital layers of the model, as described in Section 3.2.